OpenStudy (anonymous):

The sides of a triangle measure 14.9 centimeters, 23.8 centimeters, and 36.9 centimeters. Find the measure of the angle with the least measure.

OpenStudy (raden):

the least measure is the angle infront the least side, use the cosine rule : cos x = [(14.9)^2 + (36.9)^2 - (23.8)^2]/(2*14.9*36.9)

OpenStudy (raden):

simplify right side, then use arc(cos) to get x and dont forget use calculator :P

OpenStudy (anonymous):

i get domain error when i plug it in

OpenStudy (agent0smith):

^ you may have made a mistake in your calculation. \[\cos x = \frac{ (14.9)^2 + (36.9)^2 - (23.8)^2 }{(2*14.9*36.9) } = 0.925\]